### v2

# Exploring Network-Knowledge Graph Duality: A Case Study in Agentic Supply Chain Risk Analysis

Abstract.
Large Language Models (LLMs) struggle with the complex, multi-modal, and network-native data underlying financial risk. Standard Retrieval-Augmented Generation (RAG) oversimplifies relationships, while specialist models are costly and static. We address this gap with an LLM-centric agent framework for supply chain risk analysis. Our core contribution is to exploit the inherent duality between networks and knowledge graphs (KG). We treat the supply chain network as a KG, allowing us to use structural network science principles for retrieval. A graph traverser, guided by network centrality scores, efficiently extracts the most economically salient risk paths. An agentic architecture orchestrates this graph retrieval alongside data from numerical factor tables and news streams. Crucially, it employs novel “context shells” — descriptive templates that embed raw figures in natural language — to make quantitative data fully intelligible to the LLM. This lightweight approach enables the model to generate concise, explainable, and context-rich risk narratives in real-time without costly fine-tuning or a dedicated graph database.
copyright: none
1. Introduction
Large language models (LLMs) are increasingly embedded in critical, high-stakes decision processes, from medical diagnostic triage (dedhia2025bupsi) to financial advice(takayanagi2025generative). However, most domain deployments still rely on fine‑tuning a specialist model on a hand‑curated corpus or on straightforward retrieval augmented generation (RAG) that proxies relationships with vector distance. Both paths leave value on the table for supply chain risk analysis: the former is costly to update with current events and freezes knowledge at training time, while the latter ignores the rich network semantics hidden in structured data. In addition, they typically operate on a single modality, despite the fact that domain‑specific tasks require seamless reasoning over text, tables, graphs, and time‑series data.
We address this gap with an LLM‑centric system that uses network science principles to uncover and expose semantically meaningful supply‑chain paths, for example, Apple $\leftrightarrow$ Smartphones $\leftrightarrow$ Integrated Circuits. These distilled sub-graphs are fed into the LLM’s context at inference time. The paths are first retrieved by traversing the supply chain network from seed nodes that are semantically similar to key entities in a user’s input. Using structural centrality metrics to determine traversal distance, economically relevant sub‑networks are retrieved and then formatted to expose semantic significance to the LLM. Beyond graph paths, our framework ingests numerical factor tables and curated news snippets, fusing these modalities into a single prompt so the model can weigh quantitative signals alongside narrative context. This turns the graph into an “explainable retrieval engine” rather than a static data source. By importing only the highest‑salience paths, we give the LLM a concise, interpretable scaffold on which to reason, while freeing it to generate natural‑language risk narratives.
Why a network lens? Vector similarity treats every fact as an isolated point; supply‑chain risk lives in the links. By casting the KG as a graph $G=(V,E)$ whose edges denote economic relations (Produces, Has Input, Manufactured In), we use network analysis to avoid over-flooding the LLM’s context window. Therefore, each retrieved path is (i) interpretable—a concise Company → Product → Location narrative the agent can inspect—and (ii) actionable: the same traversal provides quantitative signals such as edge weights for the revenue generated by a particular product for a given company. The result is a prompt that carries built‑in economic meaning, enabling the LLM to explain hidden dependencies rather than guess relationships from token proximity.
Our approach is built on a foundational insight: the duality of the supply chain network and the knowledge graph. A network is a set of nodes and edges, whereas a knowledge graph is a set of entities and semantic relationships. For supply chains, these are one and the same. The economic edge (Company A)-[PRODUCES]-¿(Product B) is both a structural link in a network and a semantic triple in a knowledge graph. This duality allows us to reframe the complex problem of KG traversal into an efficient network-science problem, using well-established centrality metrics to identify salient paths for LLM reasoning.
Contributions. We pose four main advances:
  1. (1)
    Network‑science path discovery that extracts relevant sub-graphs.
  2. (2)
    Knowledge-graph semantic encoding transforms a network into inference‑time prompts exposing built‑in economic meaning.
  3. (3)
    Agent‑orchestrated, multi‑modal retrieval loop where a triage agent selects between graph traversal, factor data, and news tools prior to synthesis.
  4. (4)
    Context shells for numerical data that wrap each figure in descriptive language, letting the LLM reason over context-rich quantitative risk metrics.
Paper organization. Section 2 reviews graph-aware LLM curricula, GraphRAG, and graph database traversal. Section 3 sketches the overall agentic architecture and loop. Section 4 describes the three data channels: risk factors, curated news, and synthetic supply chain graph. We also describe the tools that retrieve them, while Section 5 zooms in on the factor context-shell template. Section 6 explains how retrieved evidence is merged into the prompt and the rank-then-traverse algorithm that extracts salient supply chain paths. A live dialogue in Section 7 shows the retrieval chain in action, Section 8 concludes, and the Appendix A provides full relationship tables and the complete KG diagram.
2. Background and Related Work
Bottom-Up KG curricula.
Recent work on bottom‑up domain‑specific super‑intelligence (BDSI) shows how multi‑hop KG paths can be verbalized into 24,000 reasoning tasks that supervise a 32 B‑parameter model, yielding state‑of‑the‑art scores on ICD‑Bench (dedhia2025bupsi). Their curriculum transforms the graph structure into training data, pushing the model to compose primitives into higher‑order concepts.
Graph‑aware retrieval.
GraphRAG (edge2024graphrag) first extracts an entity knowledge graph from a document corpus, then offline generates hierarchical community summaries using a Leiden‑based clustering; at query time, it assembles partial answers from those summaries and fuses them into a global response. This excels on corpus‑wide “sense‑making” questions but its pre‑processing and multi‑stage summarization pipeline are compute‑intensive. Lighter variants such as Neural‑KB, GNN‑RAG(mavromatis2024gnn) and temporal‑aware RAG(zhu2025fintmmbench) likewise keep the base LM frozen while injecting graph structure to boost recall and freshness. For real-time applications such as finance, retraining such a model daily is even less feasible.
KG traversal in existing engines.
Platforms such as Neo4j,(neo4j_cypher_manual) and LangChain’s Graph Retriever,(bratanic2024langchain) execute a pattern like
MATCH (a)-[:SUPPLIES*1..3]->(b)
by expanding a frontier of node IDs from a seed vertex, applying the edge labels, hop range, property filters, and optional top-$k$ limit that the user specifies. Because the traversal runs inside a live Neo4j server, whose page cache keeps hot neighborhoods in memory while the full adjacency list stays on disk, results arrive in milliseconds. Nevertheless, they do require a running graph service with enough RAM to cache frequently accessed portions of the graph. Our method retains the same query vocabulary, yet materializes only the path segments needed at inference, avoiding the operational overhead of a dedicated graph database while still supporting fast, parameter-controlled look-ups.
Positioning of this paper.
Our proposed agentic framework bridges the gap: it agrees with the insight in (dedhia2025bupsi) that KG paths encode domain reasoning, but applies that insight at inference time with lightweight network traversal, avoiding the cost and staleness of specialist fine‑tuning while remaining faster than on‑the‑fly graph construction methods like GraphRAG.
3. System Overview
Upon initiating a chat, the user supplies our system with the portfolio constituents and their weights. We make sure that the LLM holds this at the top of the context window to always be able to refer to it, since LLMs are known to get “lost in the middle” and not pay attention to data in the middle of the context window(liu2023lostmiddlelanguagemodels).
Figure 1 gives a bird’s‑eye view of our framework. A user (typically a CIO) poses a query; two lightweight agents decide what to retrieve and how to format before a ‘frozen’ LLM generates the final answer. Here and in what follows, by ‘frozen’ we mean a pre-trained LLM that is not being fine-tuned as new data arrive.
Figure 1. End‑to‑end pipeline. The Triage Agent decides whether the query can be answered from memory. If not, a Rerouting Agent selects one or more tools (Factors, News, Supply‑Chain KG), whose outputs are stored in a temporary database and injected into the LLM prompt.
At the start of every turn, the user’s message is appended to the ongoing dialogue and sent, together with a snapshot of the portfolio, to a lightweight Triage Agent. When that agent finds the answer in memory, it returns it straight away; otherwise it flags the query for augmentation and delegates to a downstream Rerouting Agent that decides which retrieval tool (factor exposures, curated news or a supply-chain walk) to invoke and with what parameters.
The selected tools tap separate FAISS indices(douze2024faiss), one per modality, so semantic relationships inside each data type remain intact and compound queries [^1] A request is termed a compound query when it references two or more distinct sub-topics or entities. Embedding the full sentence typically positions its vector between the relevant clusters in space, so a $k$-nearest search can miss evidence dispersed across those clusters. For example, in the query “factor exposures and cobalt” there may be no single source covering both topics together, while each topic is well covered individually at a significant vector distance from the other; $k$-nearest retrieval therefore fails to return comprehensive coverage. can be resolved cleanly across stores. Each tool call is executed through the OpenAI function-calling interface; the language model emits a JSON stub, the back-end runs the call, and the retrieved snippets flow into a transient database before being re-injected into the prompt. Finally, a frozen GPT-4o model synthesizes those multi-modal fragments with the original question, streams an Augmented Response to the UI, and stores the turn in memory, readying the loop for whatever the user asks next. Responses are fast, as the tool calling system is a series of quick API calls.
This modular design balances interpretability (the network paths are passed verbatim), flexibility (new tools can be dropped in without re‑training), and efficiency (no fine‑tuning loop). The next section details the construction of each data mode.
4. Data Modes And Tools
Our system operates on three external data channels that complement the conversation context and the user’s portfolio. Each is stored in its own FAISS index and is retrieved through a dedicated tool.
Multi‑Asset‑Class (MAC) factors.
For every security in the user’s portfolio, we ingest MSCI MAC(shepard2020msci) factor scores and the accompanying methodology text. Before embedding, each record is wrapped in a context shell (see section 5): a short paragraph that “tokenizes” the numerical z‑scores within natural‑language sentences so the LLM can reason about their meaning rather than seeing them as opaque numbers. The resulting embeddings power the get_factors tool.
Curated news.
Two news streams feed the system. Macro articles on long‑horizon risks (demographics, climate, geopolitics) and stock‑specific news comes from the LexisNexis archive, restricted to the single trading day. All articles are chunked to roughly a page, embedded once, and served by the get_news tool; the vector store preserves outlet and timestamp metadata to support recency filters.
Supply‑chain knowledge graph.
The final mode is a multi‑entity knowledge graph that links Company, Product, Input Product, Input to Input Product, Industry, and Location nodes via relations such as Produces, Has Input, Manufactured In and so forth (the full list of edge types are given in LABEL:tab:entity_rships_kg in A. Although illustrative rather than fully industrial, the graph is constructed with a pipeline akin to AIPNet (fetzer2024aipnet), to query ChatGPT’s intrinsic knowledge base to create a dataset of the desired form. These paths enter the prompt as restructured text within a context shell, exposing the semantic relationships between nodes on the traversed graph.
Taken together, MAC factors, curated news, and the supply‑chain graph give our system the multi‑modal context required for real‑time, portfolio risk analysis.
5. Context Shell
Domain risk analysis is based on numbers: z-scored factor exposures, portfolio weights, valuation ratios. Large language models, however, treat raw figures as opaque tokens and rarely attach contextual economic meaning to them. To retain a frozen, generic LLM while still letting it reason quantitatively, our proposed system wraps every table row from the MSCI Multi-Asset-Class (MAC) factor model in a context shell: a short paragraph that “tokenizes” each figure inside an explanatory sentence so that neighboring words endow the number with semantics. The surrounding text becomes part of the embedding; the number itself is now a first-class token that via attention relationships with pertinent context becomes easily interpretable. A related but different idea has been explored in (anthropic2024contextual).
Illustrative shell. Below is an excerpt from a shell; the highlighted values are place holders for numerical figures inserted for every security whose factor exposures we have:
The position in the portfolio is associated with the security [‘Security Name’] represented by the ticker [‘Ticker’]. This position constitutes [‘Weight’]% of the total portfolio. Each of the following factors is given a z-score (mean 0, sd 1) for this equity relative to all other equities. [‘Security Name’] Equity Beta: [‘Equity Beta’]
Description: captures market risk beyond the baseline Market factor.
When High: portfolio tilts toward high-beta stocks, amplifying risk.
When Low: portfolio tilts toward low-beta stocks, partially offsetting risk.
[‘Security Name’] Book-to-Price: [‘Book-to-Price’]
Description: book value divided by market capitalization.
When High: stock may be undervalued or distressed.
When Low: stock may be overvalued or considered a growth stock.
When the shell is embedded and stored in a dedicated FAISS index, both the verbal context and the numeric tokens contribute to semantic weighting. At inference time, a get_factors tool fetches the most similar shells for the query; the LLM then attends to what the numbers mean rather than merely copying them, enabling factor-aware narratives without any task-specific fine-tuning or database querying, giving the LLM a lightweight yet principled bridge between structured factor data and natural language reasoning.
6. Network‑Science Path Discovery
Here, we operationalize the network-KG duality. Instead of complex queries over a knowledge graph database, we use network traversal algorithms to retrieve economically significant sub-graphs.
When a user query arrives, the Triage Agent inspects both the conversation history and the current portfolio, deciding whether the answer is already in memory or whether fresh data are needed. For queries that could benefit from the supply chain context, it delegates to a graph-traverser tool chosen by the downstream Rerouting Agent.
The traverser first extracts every mention of the company, product, or location in the user’s text, embeds those strings with the same model used to pre-compute node embeddings for the knowledge graph, and retrieves the closest vectors; the matched vertices become seed nodes for traversal. Because each vertex is stored inside a context-shell that exposes its business metadata during embedding, similarity search is guided by both entity names and surrounding economic meaning, yielding semantically faithful starting points.
Traversal depth automatically adapts to the structural role of each seed node. We use three network centrality measures to capture different facets of structural node importance. Degree is a local score that counts immediate connections of a node, highlighting popular “hubs”. Closeness gauges how near a node lies, on average, to all others, rewarding globally well-connected vertices, while betweenness tallies how often a node sits on shortest paths, identifying nodes that link otherwise distant regions of the graph. These unweighted scores are utilized in this prototype as an efficient first-pass filter to identify structurally important nodes (e.g., major hubs or bottlenecks for information flow), which often proxy for economic relevance in large networks. The explicit economic weights are only applied in the subsequent semantic encoding step.
Averaging the three yields a single salience value that reflects both the influence of a node on its neighborhood and its role in the global network structure, giving the traversal algorithm a balanced way to decide how far to expand from each seed node. For all nodes, this centrality metric is pre-computed. As seen when contrasting the one hop network in Figure 2 and the two hop network in Figure 3, the number of nodes retrieved increases exponentially with each additional hop.
Figure 2. Apple’s network with immediately adjacent Product nodes.
Figure 3. Two-hop supply-chain network for Apple.
In practice supply chain networks follow a scale-free, power-law degree distribution (mungo2023reconstruct): hubs such as Electronic Integrated Circuits in the middle of Figure 3 in the Apple network expose a wealth of information in just one hop, whereas peripheral nodes require additional traversal hops for useful coverage.
The subtree gathered from this traverse step is then distilled into short narrative paths. A network, strictly defined, is a collection of connected objects. Each object has minimal metadata, such as an identifier, and edges may have weights. Underneath the hood, in its most basic form, this is what a supply chain is.
For example, if traversal identifies weighted path as Apple Inc. $\rightarrow$ Desktop Computers $\rightarrow$ Integrated Circuits $\rightarrow$ Shanghai, China with the corresponding edge weights as $10\%,19\%,13\%$ respectively, our system verbalizes this structured data into an intuitive narrative. By treating the network as a knowledge graph, we leverage its inherent semantic structure, bringing it closer to natural language. Add node types and edge types that can be encoded into context shells, and transform this information into the following:
Apple generates 10% of its revenue from selling Desktop Computers, which spends 19% of its production budget on Integrated Circuits, 13% of which are produced in Shanghai, China.
The semantic text surrounding the injected entities and data is predefined by the edge type between two types of nodes. The complete set of edge types and metadata is provided in Appendix A.
Each edge type corresponds to a semantically meaningful edge phrase, so the language model receives a compact but economically meaningful scaffold on which to reason. These supply chain snippets let AI-CRO explain why a node matters, linking downstream cash flows to upstream inputs and vice versa depending on the CIO’s query.
7. Sample Chat
The dialogue below illustrates how AI-CRO chains triage, tool calls, and multi-modal retrieval in real time. The CIO manages a portfolio of the top-50 constituents of the MSCI USA Index, where Apple is the largest weight, and Tesla a significant name in electric vehicles (EVs).
User: I read about problems in the DRC with coltan.
System: Issues in the Democratic Republic of Congo (DRC) involving coltan, a key source of tantalum for electronics, overlap with risks of cobalt supply that affect battery producers. Apple and Tesla rely on both minerals, so disruptions in eastern DRC represent potential operational and reputational vulnerabilities for the portfolio. Reference: Supply-chain paths for coltan.
Our system mapped “coltan” to a product node, located the upstream path into Apple’s and Tesla’s battery lines, and surfaced a portfolio-level risk narrative.
User: What has been in the news on this?
Console (log): Tool called: get_news
Query embedding: “recent news on coltan & cobalt supply-chain issues in the DRC”
Retrieved: News article, pages 3, 4 and 1
System: Recent coverage documents coltan mined by the M23 militia in eastern DRC, smuggled via Rwanda into global supply chains. This makes it harder for Apple, Tesla, and peers to guarantee conflict-free sourcing, heightening both supply-disruption and ESG-reputation risk. Reference: News Article.
The Rerouting Agent embedded a semantic description, retrieved the top-three matching articles, and the LLM wove their content into the conversation.
User: Can you walk me through the various ways this could hurt Apple?
System: (i) Supply-chain delays — shortages of tantalum capacitors could postpone iPhone and Mac launches. (ii) Reputational risk — perceived reliance on conflict minerals may erode brand trust. (iii) Margin pressure — higher input costs compress hardware margins and may force price rises that dampen demand. Reference: Apple supply-chain paths.
Across three turns, our system triaged queries, selected tools, retrieved factor-relevant evidence, and produced narrative answers that integrate operational, ethical, and financial dimensions, without any bespoke fine-tuning.
8. Conclusion and Next Steps
Summary. Our proposed framework demonstrates that a frozen general purpose LLM can operate as a domain expert when paired with three carefully engineered data modes. MSCI MAC factor models, curated news, and a synthetic supply chain knowledge graph, along with an agent loop that decides when and how to pull each modality. By turning high-salience graph paths into natural-language snippets and wrapping numeric factors in context shells, the system surfaces hidden dependencies and quantitative signals that ordinary vector-only RAG pipelines overlook, all while running in real time without specialist fine-tuning.
Although our system demonstrates a powerful new approach, we acknowledge several limitations that point to avenues for future work. First, the supply chain graph was constructed using an LLM pipeline and is inherently limited in scope. While effective for this case study, it may contain inaccuracies or omissions. Future work should focus on validating and enriching this graph with structured trade data from sources such as customs declarations or bill of landing records.
Second, our path discovery relies on unweighted topological centrality scores to guide the traversal. While this effectively identifies structurally important paths, it can overlook economically critical paths involving peripheral nodes with a high financial weight. Future work will investigate integrating edge weights directly into the traversal algorithm (e.g., using weighted shortest path algorithms or value-weighted centrality measures) to improve economic fidelity.
Finally, the system currently uses static edge weights. A key next step is to integrate real-time financial data to dynamically update these weights, reflecting changing revenue dependencies or production costs.
Appendix A Knowledge-Graph Details
Table 1. Relationships by Entity Type for the knowledge graph.
Entity Type Possible Relationships Entity Metadata
[…]
Ticker
Total Revenue
[…]
HS Code
Total Revenue Share
Production Cost Percentage
Industry Includes $\rightarrow Product$
[…]
NAICS Code
[…]
Latitude
Production Share
Figure 4. Knowledge-graph relationship diagram.